# A New Search Pipeline for Short Gamma Ray Bursts in Fermi/GBM Data - A 50% Increase in the Number of Detections

### v1

@@ 1 Introduction
Ever since the very first serendipitous detection of Gamma-ray Bursts (GRBs) in the 1970s (Klebesadel et al. (1973)), a few thousand events have been observed (Levan (2018)). These are brief, energetic bursts of electromagnetic (EM) radiation with energies in the gamma-ray band - typically from tens of KeV to one MeV - and durations lasting from a fraction of a second to a few hundred seconds. The growing number of GRB detections, and the development of instruments with better spectral resolution, have allowed for the identification of two distinct classes based on duration and spectral hardness (Kouveliotou et al. (1993)). These classes have been further analyzed and refined with enhanced spectral data from detectors and improved statistics from larger samples (Von Kienlin et al. (2020)). In addition, the increased number of GRBs, including their localizations, allows for better characterization of their rates, hosts, and progenitors (Berger (2014); Blanchard et al. (2016); Fong et al. (2009); Fong & Berger (2013)). It is now believed that long/soft GRBs arise from the collapse of a massive star (Woosley (1993); Galama et al. (1998); Kulkarni et al. (1998); Lunnan et al. (2014)), which is based on the GRB-supernova association of tens of events (Cano et al. (2017)). For short/hard bursts, the progenitor was predicted to be a binary neutron star merger (Paczynski (1986); Goodman (1986); Eichler et al. (1989)) which was also expected to be the site of heavy element production - a so-called kilonova (Li & Paczyński (1998); Kulkarni (2005), also see Metzger (2020) for a review). This hypothesis was confirmed observationally by the joint observation of the gravitational wave event GW170817 (Abbott et al. (2017a)), an associated short GRB (Goldstein et al. (2017); Savchenko et al. (2017)), and a kilonova from the same source (Smartt et al. (2017); Tanvir et al. (2017)). This is the only GW signal event to date with a confirmed EM counterpart – the joint event was remarkably successful in constraining the central engine of GRBs in ways that would be impossible using EM signals alone (Abbott et al. (2017b)), but it also raised new questions that remain unanswered. For example, the luminosity of GRB170817A was several orders of magnitude lower than that of any previously observed GRB (Abbott et al. (2017b)). Additionally, the EM signal was detected approximately 1.7 seconds after the GW signal ended, a timing offset that remains a topic of debate (Zhang (2019); Salafia et al. (2018); Gottlieb et al. (2018)). These scientific questions underscore the importance of multimessenger astronomy, and specifically motivate investigations into maximizing our capability to detect GRBs both alone and in coincidence/follow-up with GWs.
@@
Today, the most sensitive all-sky GRB detector is the Gamma-Ray Burst Monitor (GBM; Meegan et al. (2009)) on board the Fermi satellite. GBM consists of 14 scintillation detectors, 12 of which are Sodium Iodide (NaI) crystals covering an energy range of $8-1000$ KeV, and two Bismuth Germanate (BGO) crystals covering the higher energy range of $200$ KeV - $40$ MeV. At any given time, this set of detectors is sensitive to the entire sky unocculted by the Earth, corresponding to a solid angle of approximately 8 steradians. Given that gravitational-wave detectors also operate with all-sky coverage, GBM is well-suited for finding coincident GRB–GW detections.
@@ 3.1 Template Banks
Figure 2: Comparison of templates generated with different software at different times. We matched every template in the coarse bank, generated using the response generator gbm_drm_gen at met=654393605.0 sec (2021-09-27 00:00:00), with the same spectral parameters and sky positions but generated using GBM’s response generator at the specified dates which where randomly chosen. On 2018-01-17 at 16:51:09, Fermi was in an offset-pointing position due to a Target of Opportunity (ToO) observation (Obs ID 112). The figure indicates that the SNR loss is small and is bounded by less than 10%. Even during a ToO, when the spacecraft was at a different angle with respect to Earth than usual, the maximal mismatch was only 15%.
@@
$F_{E}=A\begin{cases}\left(\frac{E}{100\mathrm{KeV}}\right)^{\alpha}\exp\left(-\frac{E}{E_{0}}\right),E<(\alpha-\beta)E_{0},\\ \left[\frac{(\alpha-\beta)E_{0}}{100\mathrm{KeV}}\right]^{\alpha-\beta}\exp(\beta-\alpha)\left(\frac{E}{100\mathrm{KeV}}\right)^{\beta},\\ \qquad\qquad\qquad\qquad\qquad E\geq(\alpha-\beta)E_{0},\end{cases}$ (2)
@@
Figure 4: Template space coverage checks. Left: Match of each template in the reference bank with all the templates in the coarse bank, keeping the maximal value. The y-axis is the cumulative fraction of templates with a maximal match larger than the number specified by the x-axis. The reference bank is a larger bank, containing $5\times 10^{4}$ templates. When using Poisson statistics, only a small fraction, $\sim 0.02$, has a match less than $0.95$. Right: Comparison of the GBM template bank and our template bank. One can see that our $\sim 500$ templates are able to cover more than $90\%$ of the parameter space covered by the GBM search bank containing $3$ spectral shapes for each of the $10^{4}$ sky positions required for $\sim 4$ deg2 resolution.
@@ 3.2 Data pre-processing
The data format we use in our analysis is TTE, provided by the GBM team, and is publicly available on the HEASARC FTP server. This data format provides the highest energy resolution, as well as temporal resolution. We also use some of the utilities of GBM data tools python package (Goldstein et al. (2022)) provided by the GBM team, such as binning the TTE data, plotting sky maps and converting angles to and from the spacecraft coordinate system. Our basic analysis unit is one TTE file, consisting of one hour of photon counts per detector. We bin the data to either 0.01 seconds or 0.001 seconds, depending on the GRB duration we search for. The TTE file also holds information about the detector’s Good Time Intervals (GTI). Since we analyze all the detectors coherently, we use the intersection of all the detectors’ GTIs as the GTI for the data segment. To avoid edge effects, we exclude the first and last 15 seconds around any GTI. We also exclude data segments in which at least one detector is off for more than several tens of milliseconds. Another bad data segment we filter out relates to extremely high count rates that cause pulse pile-up and, therefore, non-reliable detector behavior (similar to what happened with GRB221009A). This effect occasionally occurs due to strong SFs and is not always included in the GTI reported. Lastly, in the time-slided data and simulated data, we exclude all catalog events (e.g., SFs and TGFs), whereas in the original data, we only exclude catalog SFs.
@@ 3.3.2 Local Background Count Estimation
For the constant background term in eq. 5, we take the average of the background estimator 7 in short windows (generally $2-4$ min), effectively splitting the analysis into independent segments $\hat{b}_{n}=1/T\sum_{t}\hat{b}_{t,n}$. The estimation error in the log term of the statistic is less sensitive to estimation errors.
@@ 3.4.5 Trigger Direction
[…]
@@ 4 Results
This distribution includes all detected triggers, encompassing SGRs, SFs, and TGFs. By applying selection criteria based on trigger properties—such as location statistics (see sec. 3.4.5) or peak energy—it is possible to exclude non-GRB events. However, this filtering approach is relatively rudimentary and may remove genuine GRBs. A more robust classification framework is therefore necessary. The development of such a scheme, along with results from the complete GBM dataset, will be presented in a separate paper. Additional discussion of the classification method and its implications can be found in sec. 5. Fig. 12 presents a scatter plot of peak energy versus duration, with points color-coded by $\mathcal{B}_{\oplus}$, illustrating the effects of the applied cuts. Distinct clustering is observed in the short-duration, high-energy region (upper left), as well as in the longer-duration, lower-energy region (lower right). However, the plot clearly demonstrates that simple cuts are insufficient for definitive classification. Within the dashed region, where sGRBs are expected, events with high Earth association are still present, and conversely, cataloged GRBs are found outside this region. Furthermore, some events within the sGRB region may instead originate from SGRs, highlighting the limitations of this approach.
@@
Figure 14: Maximum likelihood sky distribution of triggers with $p_{\text{astro}}\geq 1$. Left: All triggers with $p_{\text{astro}}\geq 1$ and $\mathcal{B_{\oplus}},\mathcal{B_{\odot}}<1$. One can see the over-density of triggers close to the galactic center. These triggers are relatively soft, with $E_{\text{peak}}<40$ KeV. Right: Similar to the left panel but now applying a peak energy cuts, keeping only triggers with $40\;\text{KeV}<E_{\text{peak}}<2000\;\text{KeV}$.
@@ 5 Conclusions and Discussion
Our enhanced detection capabilities, which yield a $50\%$ increase in the number of sGRB, can play an important role in the search for coincident detections of sGRBs and GWs from LIGO observing runs. Conducting targeted GW searches around the newly identified sGRBs, while incorporating their sky localization posteriors as priors, has the potential to increase the joint detection volume and improve the sensitivity of multi-messenger searches. This will be further enhanced by incorporating the larger sample of sGRB candidates with $10.5<p_{\text{astro}}<$ 1, which will be available upon the application of our pipeline over the entire GBM dataset.
@@ Acknowledgments
This research was supported by grant no 2022136 from the United States - Israel Binational Science Foundation (BSF), Jerusalem, Israel BZ is supported by a research grant from the Willner Family Leadership Institute for the Weizmann Institute of Science.
@@ Appendix E Sky posterior
$=\int\text{d}\xi\;\exp{\sum_{n}\left[d_{n}\log\left(1+\frac{AT_{n}(\Omega,\xi)}{b_{n}}\right)-AT_{n}(\Omega,\xi)\right]}.$ ) . (E4)
@@ Appendix F Bayesian Model Selection
[…]
@@
[…]
@@ Appendix I Contaminant Removal Tests Histograms
GBM Pipeline

### v2

@@ 1 Introduction
Ever since the very first serendipitous detection of Gamma-ray Bursts (GRBs) in the 1970s (Klebesadel et al. (1973)), a few thousand events have been observed (Levan (2018)). These are brief, energetic bursts of electromagnetic (EM) radiation with energies in the gamma-ray band - typically from tens of KeV to one MeV - and durations lasting from a fraction of a second to a few hundred seconds (Kouveliotou et al. (1993); Paciesas et al. (1999); Von Kienlin et al. (2020); Lien et al. (2016)). The growing number of GRB detections, and the development of instruments with better spectral resolution, have allowed for the identification of two distinct classes based on duration and spectral hardness (Kouveliotou et al. (1993)). These classes have been further analyzed and refined with enhanced spectral data from detectors and improved statistics from larger samples (Von Kienlin et al. (2020)). In addition, the increased number of GRBs, including their localizations, allows for better characterization of their rates, hosts, and progenitors (Berger (2014); Blanchard et al. (2016); Fong et al. (2009); Fong & Berger (2013)). It is now believed that long/soft GRBs arise from the collapse of a massive star (Woosley (1993); Galama et al. (1998); Kulkarni et al. (1998); Lunnan et al. (2014)), which is based on the GRB-supernova association of tens of events (Cano et al. (2017)). For short/hard bursts, the progenitor was predicted to be a binary neutron star merger (Paczynski (1986); Goodman (1986); Eichler et al. (1989)) which was also expected to be the site of heavy element production - a so-called kilonova (Li & Paczyński (1998); Kulkarni (2005), also see Metzger (2020) for a review). This hypothesis was confirmed observationally by the joint observation of the gravitational wave event GW170817 (Abbott et al. (2017a)), an associated short GRB (Goldstein et al. (2017); Savchenko et al. (2017)), and a kilonova from the same source (Smartt et al. (2017); Tanvir et al. (2017)). This is the only GW signal event to date with a confirmed EM counterpart – the joint event was remarkably successful in constraining the central engine of GRBs in ways that would be impossible using EM signals alone (Abbott et al. (2017b)), but it also raised new questions that remain unanswered. For example, the luminosity of GRB170817A was several orders of magnitude lower than that of any previously observed GRB (Abbott et al. (2017b)). Additionally, the EM signal was detected approximately 1.7 seconds after the GW signal ended, a timing offset that remains a topic of debate (Zhang (2019); Salafia et al. (2018); Gottlieb et al. (2018)). These scientific questions underscore the importance of multimessenger astronomy, and specifically motivate investigations into maximizing our capability to detect GRBs both alone and in coincidence/follow-up with GWs.
@@
Today, the most sensitive all-sky GRB detector is the Gamma-Ray Burst Monitor (GBM; Meegan et al. (2009)) on board the Fermi satellite. GBM consists of 14 scintillation detectors, 12 of which are Sodium Iodide (NaI) crystals covering an energy range of 810008-1000 KeV, and two Bismuth Germanate (BGO) crystals covering the higher energy range of $200$ KeV - $40$ MeV. At any given time, this set of detectors is sensitive to the entire sky unocculted by the Earth, corresponding to a solid angle of approximately 8 steradians. Given that gravitational-wave detectors also operate with all-sky coverage, GBM is well-suited for finding coincident GRB–GW detections.
@@ 3.1 Template Banks
Figure 2: Comparison of templates generated with different software at different times. We matched (as in Eq. 3.1) every template in the coarse bank, generated using the response generator gbm_drm_gen at met=654393605.0 sec (2021-09-27 00:00:$0$), with templates having the same spectral parameters and sky positions but generated using GBM’s response generator at the specified dates, which were randomly chosen. One can see that the maximal mismatch is only 15%, and that for the most part, the time in which the response was generated does not affect the template significantly. On 2018-01-17 at 16:51:09 (brown line), Fermi was in an offset-pointing position due to a Target of Opportunity (ToO) observation (Obs ID 100601-1-1), which resulted in a slightly higher mismatch. A full table describing the observing modes during these times is presented in Appendix J.
@@
[…]
@@
Figure 4: Template space coverage checks. Match of each template in the reference template bank and each template in the GBM-like bank with all the templates in the coarse bank, keeping the maximal value. The y-axis is the cumulative fraction of templates with a maximal match larger than the number specified by the x-axis. The reference bank is a larger bank, containing $5\times 10^{4}$ templates. We see that only a small fraction, $\sim 0.02$, has a match less than $0.95$. Compared with the GBM-like templates, our $\sim 500$ templates cover more than $90\%$ of the parameter space covered by the GBM search bank containing 33 spectral shapes for each of the $10^{4}$ sky positions required for $\sim 4$ deg2 resolution.
@@ 3.2 Data pre-processing
The data format we use in our analysis is TTE, provided by the GBM team, and is publicly available on the HEASARC FTP server. This data format provides the highest energy resolution, as well as temporal resolution. We also use some of the utilities of GBM data tools python package (Goldstein et al. (2022)) provided by the GBM team, such as binning the TTE data, plotting sky maps and converting angles to and from the spacecraft coordinate system. Our basic analysis unit is one TTE file, consisting of one hour of photon counts per detector. We bin the data to either 0.01 seconds or 0.001 seconds, depending on the GRB duration we search for. The TTE file also holds information about the detector’s Good Time Intervals (GTI). Since we analyze all the detectors coherently, we use the intersection of all the detectors’ GTIs as the GTI for the data segment. To avoid edge effects, we exclude the first and last 15 seconds around any GTI. We also exclude data segments in which at least one detector is off for more than several tens of milliseconds. Another bad data segment we filter out relates to extremely high count rates that cause pulse pile-up and, therefore, non-reliable detector behavior (similar to what happened with GRB221009A. See e.g. Lesage et al. (2023)). This effect occasionally occurs due to strong SFs and is not always included in the GTI reported. Lastly, in the time-slided data and simulated data, we exclude all catalog events (e.g., SFs and TGFs), whereas in the original data, we only exclude catalog SFs.
@@ 3.3.2 Local Background Count Estimation
For the constant background term in eq. 5, we take the average of the background estimator 7 in short windows (generally 242-4 min), effectively splitting the analysis into independent segments $\hat{b}_{n}=1/T\sum_{t}\hat{b}_{t,n}$. The estimation error in the log term of the statistic is less sensitive to estimation errors.
@@ 3.4.5 Trigger Direction
$\mathcal{B}_{\oplus}=\frac{\int d\Omega\;p(\Omega)\;I(\Omega\in\Omega_{\oplus})p(\Omega|d,\mathcal{H}_{1})}{\int d\Omega\;p(\Omega)\;I(\Omega\notin\Omega_{\odot,\oplus})p(\Omega|d,\mathcal{H}_{1})},$ (14)
@@ 4 Results
This distribution includes all detected triggers, encompassing Soft Gamma Repeaters (SGRs), SFs, and TGFs. By applying selection criteria based on trigger properties—such as location statistics (see sec. 3.4.5) or peak energy—it is possible to exclude non-GRB events. However, this filtering approach is relatively rudimentary and may remove genuine GRBs. A more robust classification framework is therefore necessary. The development of such a scheme, along with results from the complete GBM dataset, will be presented in a separate paper. Additional discussion of the classification method and its implications can be found in sec. 5. Fig. 12 presents a scatter plot of peak energy versus duration, with points color-coded by $\mathcal{B}_{\oplus}$, illustrating the effects of the applied cuts. Distinct clustering is observed in the short-duration, high-energy region (upper left), as well as in the longer-duration, lower-energy region (lower right). However, the plot clearly demonstrates that simple cuts are insufficient for definitive classification. Within the dashed region, where sGRBs are expected, events with high Earth association are still present, and conversely, cataloged GRBs are found outside this region. Furthermore, some events within the sGRB region may instead originate from SGRs, highlighting the limitations of this approach.
@@
Figure 14: Maximum likelihood sky distribution of triggers with $p_{\text{astro}}=1$. Left: All triggers with $p_{\text{astro}}=1$ and $\mathcal{B_{\oplus}},\mathcal{B_{\odot}}<1$. One can see the over-density of triggers close to the galactic center. These triggers are relatively soft, with $E_{\text{peak}}<40$ KeV. Right: Similar to the left panel but now applying a peak energy cuts, keeping only triggers with $40\;\text{KeV}<E_{\text{peak}}<2000\;\text{KeV}$.
@@ 5 Conclusions and Discussion
Our enhanced detection capabilities, which yield a $50\%$ increase in the number of sGRB, can play an important role in the search for coincident detections of sGRBs and GWs from LIGO observing runs. Conducting targeted GW searches around the newly identified sGRBs, while incorporating their sky localization posteriors as priors, has the potential to increase the joint detection volume and improve the sensitivity of multi-messenger searches. This will be further enhanced by incorporating the larger sample of sGRB candidates with $0.5<p_{\text{astro}}<1$, which will be available upon the application of our pipeline over the entire GBM dataset.
@@ Acknowledgments
This research was supported by grant no 2022136 from the United States - Israel Binational Science Foundation (BSF), Jerusalem, Israel. BZ is supported by a research grant from the Willner Family Leadership Institute for the Weizmann Institute of Science. TV additionally acknowledges support from NSF grants 2012086 and 2309360, the Alfred P. Sloan Foundation through grant number FG-2023-20470, and the Hellman Family Faculty Fellowship during the time this work was done.
@@ Appendix E Sky posterior
[…]
@@ Appendix F Bayesian Model Selection
$\mathcal{B}_{\oplus}=\frac{P(\mathcal{H}_{\oplus}|d)}{P(\mathcal{H}_{1}|d)}=\frac{\pi(\mathcal{H}_{\oplus})}{\pi(\mathcal{H}_{1})}\frac{\int d\Omega d\xi\;p(\Omega)p(\xi)L(d|\mathcal{H}_{\oplus};\Omega,\xi)}{\int d\Omega d\xi\;p(\Omega)p(\xi)L(d|\mathcal{H}_{1};\Omega,\xi)},$ (F2)
@@
$\mathcal{B}_{\oplus}=\frac{P(\mathcal{H}_{\oplus}|d)}{P(\mathcal{H}_{1}|d)}\frac{\int d\Omega\;p(\Omega|d;\mathcal{H}_{1})\;I(\Omega\in\Omega_{\oplus})}{\int d\Omega\;p(\Omega|d;\mathcal{H}_{1})\;I(\Omega\notin\Omega_{\odot,\oplus})}.$ (F3)
@@ Appendix I Contaminant Removal Tests Histograms
Table 3: Fermi observation context for the dates in Fig. 2.
Taken from https://fermi.gsfc.nasa.gov/ssc/observations/timeline/posting/
Obs ID Start (UTC) Stop (UTC) Mode Observation Details
070100-56-1 2014-10-15 01:06:00 2014-10-16 00:00:00 Pointed Modified survey for galactic center, Pole position = 68.70, 64.40. Target position (J2000): 268.44, 24.28
080400-45-1 2016-03-31 00:10:00 2016-04-07 00:00:00 Survey 50deg 6am rock; orbit period 5718 sec; -/+ 50deg rock angle over 2 orbits
090400-66-1 2017-05-18 00:10:00 2017-05-24 01:57:00 Survey 50deg 6am rock; orbit period 5715 sec; -/+ 50deg rock angle over 2 orbits
090400-71-1 2017-06-15 00:10:00 2017-06-22 00:00:00 Survey 50deg 6am rock; orbit period 5715 sec; -/+ 50deg rock angle over 2 orbits
090400-82-1 2017-07-20 00:55:00 2017-07-27 00:00:00 Survey 50deg 6am rock; orbit period 5715 sec; -/+ 50deg rock angle over 2 orbits
100601-1-1 2018-01-16 17:57:00 2018-01-17 23:12:00 ToO Pointed Offset pointing position. Target position (J2000): 171.6, -55.5
100403-6-1 2018-04-26 00:10:00 2018-04-30 14:35:00 Survey +50 deg only rock
110403-28-1 2019-01-03 17:20:00 2019-01-07 05:21:00 Survey -50 deg only rock
150401-16-1 2022-12-15 01:27:00 2022-12-22 00:00:00 Survey +50, -50 two orbit rock, 5710s
160416-10-1 2023-10-26 01:25:00 2023-10-30 00:15:00 Survey +50, -50 two orbit rock, 5710s
140400-3-1 2021-09-23 01:03:00 2021-09-27 23:55:00 Survey 50deg; orbit period 5713 sec; -/+ 50deg rock angle over 2 orbits
footnotetext: The observation times in the table are ordered in the same order as the times used in Fig 2. The last entry shows the observation details of the time used to generate our template bank.
@@
GBM Pipeline Subthreshold/OtherffSubthreshold/Other: Indicated if another subthreshold search or other instrument has detected this candidate.